# Quantum mixing on large Schreier graphs

### v1

@@ 6.1.4 Weakly converging sequences of permutation representations
Our theorems require that a sequence of finite Schreier graphs converges in distribution to the Cayley graph, here on a free product $\Gamma=\Gamma_{1}*\cdots*\Gamma_{m}$. Such sequences indeed exist and are common. Recall that $n_{j}$ is the order of $\Gamma_{j}$ and let $q$ be the least common multiple of the $n_{j}$’s. We set $N=N_{n}=qn$. Then, the set of permutation representations $\mathrm{Hom}(\Gamma,S_{N})$ is non-empty for all $n\geq 1$. Moreover, as $n\to\infty$, it follows from [66] that a uniformly random element $\rho_{N}\in\mathrm{Hom}(\Gamma,S_{N})$ converges in distribution toward the regular representation of $\Gamma$ in probability.
@@ 6.2 Right-angled Coxeter groups
In this section, we check that Theorems 1.3-1.5-1.7 apply to a family of right-angled Coxeter groups (RACG). This is quite a specific class of groups, but our purpose it to show that our results apply to classes of finitely presented groups which are not free products. Until now the main tool for checking resolvent bounds needed for quantum ergodicity (see for instance [7]) has been to write the resolvent in a product form like (6.8), but this is specific to tree-like structures and free-product structures. For RACG we need to develop another approach. Along the way, we provide methods to check the assumption (1.15) in more general groups. We start with general ideas to verify the resolvent condition (1.15) in Section 6.2.1, then we show in Section 6.2.2 that they apply in a specific class of RACG, and finally we show in Section 6.2.3 that in a subclass, there exist convergent sequences of permutation representations. Our main result of this section is Corollary 6.13. A key step in the proof is Proposition 6.12.
@@
the set of generators. Since $m_{ii}=1$, all generators are of order $2$. It follows from [28] that all RACG have property RD.
@@
Let $\Gamma$ be a RACG with set of generators $S$, whose defining diagram is super-flexible and contains three non-commuting generators commuting with all other generators. Then in ${\rm Cay}(\Gamma,S)$, both (1.5) and (1.15) hold for $p=\mathbf{1}_{S}$ across the whole spectrum except a possibly finite number of points. Moreover there exist sequences of permutation representations of $\Gamma$ of growing dimension converging in distribution toward $\lambda$. As a consequence, Theorem 1.5 applies.
@@
According to Lemma 6.9 we only need to prove (6.11). By (2.17), there exists $C_{0}$ uniform in $g$ and $\eta$ such that
$\eta|(\Im R^{z})(e,g)|^{2}e^{c|g|}\leq C_{0}.$
Let $\varepsilon>0$. There exists $R>0$ such that $C_{0}e^{-c|g|}|g|^{C_{1}^{\prime}}\leq\varepsilon$ for any $g\in\Gamma$ with $|g|>R$. Hence for any such $g$,
$\eta|(\Im R^{z})(e,g)|^{2}|g|^{C_{1}^{\prime}}\leq C_{0}e^{-c|g|}|g|^{C_{1}^{\prime}}\leq\varepsilon.$ (6.13)
@@
We claim that the family of automorphisms $\Psi_{\phi}^{j}$ allows to construct for each $g\in\Gamma$ a sufficiently large equivalence class (i.e., a sufficiently large orbit under $\mathrm{Stab}(\Gamma)$) to prove Lemma 6.17. Consider a reduced word $w$, of length $n\geq 4|S|^{2}$, and representing an element $g\in\Gamma$. There is one letter, let us call it $s$, which appears at least $n/|S|$ times in $w$. Since $s$ is of order $2$ and $w$ is reduced, between two occurrences of $s$, there is necessarily another letter which does not commute with $s$. We call the subword between two occurrences of $s$ an “$s$-segment”. So one letter $t$ which does not commute with $s$ appears in $k\geq\frac{n}{|S|^{2}}-1$ $s$-segments. We denote by $\phi$ the automorphism provided by the definition of (super-flexible) defining diagrams (with $s,t$). There are two possibilities corresponding to the two bullet points in Definition 6.10. These possibilities are symmetric in $s$ and $t$. Notice also that $s$ appears in at least (k1)(k-1) $t$-segments. Therefore, up to exchanging the roles of $s$ and $t$, we may assume that $\phi$ fixes $s$, fixes all neighbors of $s$, exchanges $t$ and $u$ for some $u$ not commuting with $s$ nor $t$, and that $t$ appears in at least $k-1\geq\frac{n}{2|S|^{2}}$ $s$-segments.
@@
$(\Psi_{\phi}^{j_{1}}\circ\ldots\circ\Psi_{\phi}^{j_{\ell}})(w)$ (6.15)
where $0\leq\ell\leq k$ and $1\leq j_{1}<j_{2}<\ldots<j_{\ell}\leq k$ (for $\ell=0$ (6.15) is equal to $w$). This yields $2^{k}$ words. We prove below that no two of these words correspond to the same element of the group. Therefore this construction yields $2^{k}$ distinct elements in the orbit under $\mathrm{Stab}(\Gamma)$ of the group element represented by $w$. Since $k\geq\frac{n}{2|S|^{2}}$, this concludes the proof of Lemma 6.17.
So still for the same $w$, let us prove that no two words of the form (6.15) correspond to the same element of the group. Assume that $(\Psi_{\phi}^{j_{1}}\circ\ldots\circ\Psi_{\phi}^{j_{\ell}})(w)$ and $(\Psi_{\phi}^{j_{1}^{\prime}}\circ\ldots\circ\Psi_{\phi}^{j_{\ell^{\prime}}^{\prime}})(w)$ correspond to the same element in the group, and consider if it exists (for the sake of a contradiction) the least $i\leq\min(\ell,\ell^{\prime})$ for which $j_{i}\neq j_{i}^{\prime}$ (or $i=\min(\ell,\ell^{\prime})+1$ otherwise). Without loss of generality, since the $\Psi_{\phi}^{j}$ are invertible, we may assume that $i=1$, and we set $j:=j_{1}$. We write
@@
where $s$ appears exactly (j1)(j-1) times in $w_{1}$. This implies (after some simplification on the left) that
@@
for some words $w_{4}$ and $w_{5}$ of same length. The equality is meant in $\Gamma$. Since $(\Psi_{\phi}^{j_{1}}\circ\ldots\circ\Psi_{\phi}^{j_{\ell}})(w)$ and $(\Psi_{\phi}^{j_{1}^{\prime}}\circ\ldots\circ\Psi_{\phi}^{j_{\ell^{\prime}}^{\prime}})(w)$ are reduced words (due to Point 2 above), we know that both members in the above identity are reduced. Therefore, according to Point 2 in Theorem 6.18, there is a sequence of commutations which transforms $w_{2}sw_{4}$ into $\phi(w_{2})sw_{5}$. We will show that this is not possible.
@@
This proves that $(\Psi_{\phi}^{j_{1}}\circ\ldots\circ\Psi_{\phi}^{j_{\ell}})(w)$ and $(\Psi_{\phi}^{j_{1}^{\prime}}\circ\ldots\circ\Psi_{\phi}^{j_{\ell^{\prime}}^{\prime}})(w)$ correspond to different elements in the group. ∎
@@ 6.3 Lifts of a finite base graph
In this subsection, we apply Theorem 5.3 to $N$-lifts of a fixed finite base graph $H=(V,E)$ (multiple edges and loops are allowed). We assume without loss of generality that $V=\left[r\right]$ and $E$ has $d$ unoriented edges and $2d$ oriented edges. Let $\Gamma=F_{d}$ be the free group on $d$ generators. Recall that $F_{d}$ has the RD property (see [44]). As explained in Example 5.2, we associate to each free generator $g_{i}$ a directed edge $(u_{i},v_{i})$ of $H$ and to $g_{i}^{-1}$ the reversed edge $(v_{i},u_{i})$. The adjacency operator of an $N$-lift of $H$ can be written as
$\rho_{N}(1_{E})=\sum_{i=1}^{d}\left(E_{u_{i}v_{i}}\otimes\rho_{N}(g_{i})+E_{u_{i}v_{i}}^{*}\otimes\rho_{N}(g_{i})^{*}\right),$
where $\rho_{N}(g_{i})=\sigma_{i}$ is some permutation matrix on $\left[N\right]$ and $1_{E}=\sum_{i}(E_{u_{i}v_{i}}g_{i}+E_{u_{i}v_{i}}^{*}g_{i}^{-1})\in M_{r}(\mathbb{C})[\Gamma]$. It is immediate to check that $g_{i}\mapsto\rho_{N}(g_{i})$ extends uniquely to a permutation representation $\rho_{N}\in\mathrm{Hom}(F_{d},S_{N})$. For $N=1$, we recover the adjacency matrix of $H$. More generally, we consider $p=p^{*}\in M_{r}(\mathbb{C})[\Gamma]$ of the form, for some $a\in\mathbb{C}^{d}$,
$p=\sum_{g}p_{g}g=\sum_{i=1}^{d}\left(a_{i}E_{u_{i}v_{i}}g_{i}+\bar{a}_{i}E^{*}_{u_{i}v_{i}}g^{-1}_{i}\right).$ (6.16)
@@
Our main result of this section is the following:
@@
This proposition has the following corollary. Let $(\sigma_{i})_{i\in\left[d\right]}$ be independent and uniform random permutations in $S_{N}$. Let $\rho_{N}\in\mathrm{Hom}(F_{d},S_{N})$ defined by $\rho_{N}(g_{i})=\sigma_{i}$. It is proved in [15] that $\rho_{N}$ converges strongly in distribution to $\lambda$ in probability as $N\rightarrow+\infty$. Consequently, from Theorem 5.3(iii) applied to $\rho_{N}(p)$, we get:
@@
Regarding assumption (5.4), since $H$ is a finite connected graph with minimal degree $d\geq 2$ which is not a cycle, [7, Proposition 4.2] applies (as is explicitly written in [7, Section 4.3], see below). Therefore, (5.4) holds for any $I_{0}$ which is contained in the interior of $\sigma(\lambda(p))\setminus D$ where $D$ is a finite set containing a possibly finite set of atoms.
We need to check (5.5). Our first step is to rephrase (5.5) in terms of the decay of the resolvent entries of the universal covering tree of $H$. For this we partly follow [16, Appendix]. For $i\in\left[d\right]$, we set $g_{i+d}=g_{i}^{-1}$, $i^{*}=i+d$ and $(i+d)^{*}=i$. The resolvent $R^{z}=(\lambda(p)-z\text{Id})^{-1}$ evaluated at $(a,b)\in\Gamma\times\Gamma$ is an element of $M_{r}(\mathbb{C})$. Let $\lambda^{(e)}(p)$ be the restriction of $\lambda(p)$ to $\mathbb{C}^{r}\otimes\ell^{2}(\Gamma\setminus\{e\})$. For $z\in\mathbb{C}\setminus\sigma(\lambda^{(e)}(p))$, we denote by
@@
the resolvent in the subtree of $\mathrm{Cay}(F_{d},(g_{i})_{i\in\left[2d\right]})$ emanating from $g_{i}$, and we let $Z^{z}_{i}=-p_{g_{i}}\gamma^{z}_{i}\in M_{r}(\mathbb{C})$. We first prove that for any $g\in\Gamma$ written in reduced form as $g=g_{i_{\ell}}\ldots g_{i_{1}}$, with $i_{k}\in\left[2d\right]$, we have
$R^{z}(e,g)=R^{z}(e,e)Z^{z}_{i_{1}}\ldots Z^{z}_{i_{\ell}}.$ (6.17)
For $i\in[2d]$, we denote by $V_{i}\subset\Gamma$ the subset of group elements $h\in\Gamma$ written in reduced form $h=g_{j_{k}}\ldots g_{j_{2}}g_{j_{1}}$ with $j_{1}=i$ and $k\in\mathbb{N}$. A sequence $(j_{1},\ldots,j_{k})$ such that $g=g_{j_{k}}\ldots g_{j_{1}}$ determines a walk of length $k$ in $\Gamma$ starting from $e$ and ending at $g$. Since $\Gamma$ is the free group and the $g_{i}$’s are the free generators, we may decompose this walk at the last passage at $e$, $g_{i_{1}}$, $g_{i_{2}}g_{i_{1}}$, $\ldots$, $g_{i_{\ell}}\ldots g_{i_{1}}=g$. This induces a decomposition of the sequence $(j_{1},\ldots,j_{k})$ into $\ell+1$ sequences $u_{0},\ldots,u_{\ell}$ (possibly empty). For any $0\leq t\leq\ell$, the sequence $u_{t}$ is a walk from $g_{i_{t}}$ to $g_{i_{t}}$ in $V_{i_{t}}$ and $(j_{1},\ldots,j_{k})=(u_{0},i_{1},u_{1},\ldots,i_{\ell},u_{\ell})$. Writing $u_{t}$ as $u_{t}=(j_{t,1},\ldots,j_{t,n_{t}})$, we thus have (setting just here $p_{i}=p_{g_{i}}$)
$\sum_{k}\frac{\lambda(p)^{k}(e,g)}{z^{k}}=\sum_{\begin{subarray}{c}(j_{1},\ldots,j_{k})\\ g=g_{j_{k}}\ldots g_{j_{1}}\end{subarray}}\frac{p_{j_{k}}\ldots p_{j_{1}}}{z^{k}}=\sum_{u_{0},\ldots,u_{\ell}}\frac{p_{j_{\ell,n_{\ell}}}\ldots p_{j_{\ell,1}}}{z^{n_{\ell}}}\frac{p_{i_{\ell}}}{z}\ldots\frac{p_{j_{1,n_{1}}}\ldots p_{j_{1,1}}}{z^{n_{1}}}\frac{p_{i_{1}}}{z}\frac{p_{j_{0,n_{0}}}\ldots p_{j_{0,1}}}{z^{n_{0}}}$
where the sum is over all sequences $(u_{0},\ldots,u_{\ell})$ as above. This concludes the proof of (6.17) for $z$ with large enough modulus, and thus for any $z$ in the upper half-plane since both expressions in (6.17) are analytic.
@@
$\gamma_{i}^{z}=-\Bigl(z\text{Id}_{r}+\sum_{j\neq i^{*}}p_{g_{j}}\gamma_{j}^{z}p_{g_{j}}^{*}\Bigr)^{-1}.$ (6.18)
@@
Let us now specify the relations (6.17) and (6.18) to the case of $p$ of the form (6.16). In this case, $p_{g_{i}}=a_{i}E_{u_{i}v_{i}}$ for all $i\in\left[2d\right]$, where for $i\in\left[d\right]$, we have set $a_{i^{*}}=\bar{a}_{i}$ and $(u_{i^{*}},v_{i^{*}})=(v_{i},u_{i})$. In this case the relation (6.18) takes a simpler form: indeed, the matrix $p_{g_{j}}\gamma_{j}^{z}p_{g_{j}}^{*}$ is diagonal, and has only one non-vanishing coefficient, in the entry $(u_{j},u_{j})$. Let $\hat{\gamma}_{i}^{z}=(\gamma_{i}^{z})_{v_{i}v_{i}}$, then (6.18) implies the scalar equation

### v2

@@ 6.1.4 Weakly converging sequences of permutation representations
Our theorems require that a sequence of finite Schreier graphs converges in distribution to the Cayley graph, here on a free product $\Gamma=\Gamma_{1}*\cdots*\Gamma_{m}$. Such sequences indeed exist and are common. Recall that $n_{j}$ is the order of $\Gamma_{j}$ and let $q$ be the least common multiple of the $n_{j}$’s. Set $N=N_{n}=qn$. Then, the set of permutation representations $\mathrm{Hom}(\Gamma,S_{N})$ is non-empty for all $n\geq 1$. Moreover, it follows from [66] that as $n\to\infty$, a uniformly random element $\rho_{N}\in\mathrm{Hom}(\Gamma,S_{N})$ converges in distribution to the regular representation of $\Gamma$ in probability.
@@ 6.2 Right-angled Coxeter groups
In this section, we check that Theorems 1.3-1.5-1.7 apply to a family of right-angled Coxeter groups (RACG). This is quite a specific class of groups, but our purpose it to show that our results apply to classes of finitely presented groups which are not free products. Until now the main tool for checking resolvent bounds needed for quantum ergodicity (see for instance [7]) has been to write the resolvent in a product form like (6.8), but this is specific to tree-like structures and free-product structures. For RACG we need to develop another approach. Along the way, we provide methods to check the assumption (1.15) in more general groups. We start with general ideas to verify the resolvent condition (1.15) in Section 6.2.1, then we show in Section 6.2.2 that they apply in a specific class of RACG, and finally we show in Section 6.2.3 that in a subclass, there exist convergent sequences of permutation representations. Our main result in this section is Corollary 6.13. A key step in the proof is Proposition 6.12.
@@
the set of generators. Since $m_{ii}=1$, all generators are of order $2$. It follows from [28] that all RACG have the RD property.
@@
Let $\Gamma$ be a RACG with a set of generators $S$, whose defining diagram is super-flexible and contains three non-commuting generators commuting with all other generators. Then in ${\rm Cay}(\Gamma,S)$, both (1.5) and (1.15) hold for $p=\mathbf{1}_{S}$ across the whole spectrum except possibly a finite number of points. Moreover there exist sequences of permutation representations of $\Gamma$ of growing dimension converging in distribution toward $\lambda$. As a consequence, Theorems 1.3 and 1.5 apply.
@@
According to Lemma 6.9 we only need to prove (6.11). By (2.17) and Lemma 6.15, there exists $C_{0}$ uniform in $g$ and $\eta$ such that
$\eta|(\Im R^{z})(e,g)|^{2}e^{c|g|}\leq c^{-1}C_{0}.$
Let $\varepsilon>0$. There exists $R>0$ such that $C_{0}c^{-1}e^{-c|g|}|g|^{C_{1}^{\prime}}\leq\varepsilon$ for any $g\in\Gamma$ with $|g|>R$. Hence for any such $g$,
$\eta|(\Im R^{z})(e,g)|^{2}|g|^{C_{1}^{\prime}}\leq C_{0}c^{-1}e^{-c|g|}|g|^{C_{1}^{\prime}}\leq\varepsilon.$ (6.13)
@@
We claim that the family of automorphisms $\Psi_{\phi}^{j}$ allows to construct for each $g\in\Gamma$ a sufficiently large equivalence class (i.e., a sufficiently large orbit under $\mathrm{Stab}(\Gamma)$) to prove Lemma 6.17. Consider a reduced word $w$, of length $n\geq 4|S|^{2}$, and representing an element $g\in\Gamma$. There is one letter, let us call it $s$, which appears at least $n/|S|$ times in $w$. Since $s$ is of order $2$ and $w$ is reduced, between two occurrences of $s$, there is necessarily another letter which does not commute with $s$. We call the subword between two occurrences of $s$ an “$s$-segment”. So one letter $t$ which does not commute with $s$ appears in $k\geq\frac{n}{|S|^{2}}-1$ $s$-segments. We denote by $\phi$ the automorphism provided by the definition of (super-flexible) defining diagrams (with $s,t$). There are two possibilities corresponding to the two bullet points in Definition 6.10. These possibilities are symmetric in $s$ and $t$. Notice also that $s$ appears in at least (k1)(k-1) $t$-segments. Therefore, up to exchanging the roles of $s$ and $t$, we may assume that $\phi$ fixes $s$, fixes all neighbors of $s$, exchanges $t$ and $u$ for some $u$ not commuting with $s$ nor $t$, and that $t$ appears in at least $k-1\geq\frac{n}{2|S|^{2}}$ $s$-segments. We define the finite set $\mathcal{I}$ of natural numbers as follows: $j\in\mathcal{I}$ if and only if there exists an $s$-segment where $t$ appears such that $s$ occurs $j$ times in $w$ before this $s$-segment. We notice that $|\mathcal{I}|\geq k-1$ since there are at least k1k-1 $s$-segments.
@@
$(\Psi_{\phi}^{j_{\ell}}\circ\ldots\circ\Psi_{\phi}^{j_{1}})(w)$ (6.16)
where $j_{1},\ldots,j_{\ell}$ are distinct elements of $\mathcal{I}$, with $j_{1}<j_{2}<\ldots<j_{\ell}$ and $\ell\leq|\mathcal{I}|$ (for $\ell=0$ (6.16) is equal to $w$). This yields $2^{|\mathcal{I}|+1}$ words. We prove below that no two of these words correspond to the same element of the group. Therefore this construction yields $2^{|\mathcal{I}|+1}\geq 2^{k}$ distinct elements in the orbit under $\mathrm{Stab}(\Gamma)$ of the group element represented by $w$. Since $k\geq\frac{n}{2|S|^{2}}$, this concludes the proof of Lemma 6.17.
So still for the same $w$, let us prove that no two words of the form (6.16) correspond to the same element of the group. Assume that $(\Psi_{\phi}^{j_{\ell}}\circ\ldots\circ\Psi_{\phi}^{j_{1}})(w)$ and $(\Psi_{\phi}^{j_{\ell^{\prime}}^{\prime}}\circ\ldots\circ\Psi_{\phi}^{j_{1}^{\prime}})(w)$ correspond to the same element in the group, and consider if it exists (for the sake of a contradiction) the least $i\leq\min(\ell,\ell^{\prime})$ for which $j_{i}\neq j_{i}^{\prime}$ (or $i=\min(\ell,\ell^{\prime})+1$ otherwise). Without loss of generality, since the $\Psi_{\phi}^{j}$ are invertible, we may assume that $i=1$, and we set $j:=\min(j_{1},j_{1}^{\prime})$. We write
@@
where $s$ appears exactly (j1)(j-1) times in $w_{1}$, and moreover $w_{2}$ is an $s$-segment containing the letter $t$. Recalling (6.15), since $j_{1}\neq j_{1}^{\prime}$, the identity $(\Psi_{\phi}^{j_{\ell}}\circ\ldots\circ\Psi_{\phi}^{j_{1}})(w)=(\Psi_{\phi}^{j_{\ell^{\prime}}^{\prime}}\circ\ldots\circ\Psi_{\phi}^{j_{1}^{\prime}})(w)$ in $\Gamma$ implies
$w_{1}sw_{2}sw_{4}=w_{1}s\phi(w_{2})sw_{5}$
for some words $w_{4}$ and $w_{5}$ of same length (because $j<\min(j_{2},\ldots,j_{\ell},j_{2}^{\prime},\ldots,j_{\ell^{\prime}})$). Thus, after simplification on the left, we get
@@
Since $(\Psi_{\phi}^{j_{1}}\circ\ldots\circ\Psi_{\phi}^{j_{\ell}})(w)$ and $(\Psi_{\phi}^{j_{1}^{\prime}}\circ\ldots\circ\Psi_{\phi}^{j_{\ell^{\prime}}^{\prime}})(w)$ are reduced words (due to Point 2 above), we know that both members in the above identity are reduced. Therefore, according to Point 2 in Theorem 6.18, there is a sequence of commutations which transforms $w_{2}sw_{4}$ into $\phi(w_{2})sw_{5}$. We will show that this is not possible.
@@
This proves that $(\Psi_{\phi}^{j_{\ell}}\circ\ldots\circ\Psi_{\phi}^{j_{1}})(w)$ and $(\Psi_{\phi}^{j_{\ell^{\prime}}^{\prime}}\circ\ldots\circ\Psi_{\phi}^{j_{1}^{\prime}})(w)$ correspond to different elements in the group. ∎
@@ 6.3 Lifts of a finite base graph
In this subsection, we apply Theorem 5.3 to $N$-lifts of a fixed finite base graph $H=(V,E)$ (multiple edges and loops are allowed). We let $V=\left[r\right]$ and denote by $d$ the number of unoriented edges. Let $\Gamma=F_{d}$ be the free group on $d$ generators. Recall that $F_{d}$ has the RD property (see [44]). As explained in Example 5.2, we associate to each free generator $g_{i}$ a directed edge $(u_{i},v_{i})$ of $H$ and to $g_{i}^{-1}$ the reversed edge $(v_{i},u_{i})$. The adjacency operator of an $N$-lift of $H$ can be written as
$\rho_{N}(1_{E})=\sum_{i=1}^{d}\left(E_{v_{i}u_{i}}\otimes\rho_{N}(g_{i})+E_{v_{i}u_{i}}^{*}\otimes\rho_{N}(g_{i})^{*}\right),$
where $\rho_{N}(g_{i})=\sigma_{i}$ is some permutation matrix on $\left[N\right]$ and $1_{E}=\sum_{i}(E_{v_{i}u_{i}}g_{i}+E_{v_{i}u_{i}}^{*}g_{i}^{-1})\in M_{r}(\mathbb{C})[\Gamma]$. It is immediate to check that $g_{i}\mapsto\rho_{N}(g_{i})$ extends uniquely to a permutation representation $\rho_{N}\in\mathrm{Hom}(F_{d},S_{N})$. For $N=1$, we recover the adjacency matrix of $H$. More generally, we consider $p=p^{*}\in M_{r}(\mathbb{C})[\Gamma]$ of the form, for some $a\in\mathbb{C}^{d}$,
$p=\sum_{g}p_{g}g=\sum_{i=1}^{d}\left(a_{i}E_{v_{i}u_{i}}g_{i}+\bar{a}_{i}E^{*}_{v_{i}u_{i}}g^{-1}_{i}\right).$ (6.17)
@@
Our main result in this section is the following:
@@
This proposition has the following corollary. Let $(\sigma_{i})_{i\in\left[d\right]}$ be independent and uniform random permutations in $S_{N}$. Let $\rho_{N}\in\mathrm{Hom}(F_{d},S_{N})$ be defined by $\rho_{N}(g_{i})=\sigma_{i}$. It is proved in [15] that $\rho_{N}$ converges strongly in distribution to $\lambda$ in probability as $N\rightarrow+\infty$. Consequently, from Theorem 5.3(iii) applied to $\rho_{N}(p)$, we get:
@@
Regarding assumption (5.4), since $H$ is a finite connected graph with minimal degree $d\geq 2$ which is not a cycle, [7, Proposition 4.2] applies (as is explicitly written in [7, Section 4.3], see below). Therefore, (5.4) holds for any $I_{0}$ which is contained in the interior of $\sigma(\lambda(p))\setminus D$ where $D$ is a finite set possibly containing atoms. We move to (5.5).
Step 1. Our first step is to rephrase (5.5) in terms of the decay of the resolvent entries of the universal covering tree of $H$, cf. (6.24) below. For this we partly follow [16, Appendix]. For $i\in\left[d\right]$, we set $g_{i+d}=g_{i}^{-1}$, $i^{*}=i+d$ and $(i+d)^{*}=i$. The resolvent $R^{z}=(\lambda(p)-z\text{Id})^{-1}$ evaluated at $(a,b)\in\Gamma\times\Gamma$ is an element of $M_{r}(\mathbb{C})$. Let $\lambda^{(e)}(p)$ be the restriction of $\lambda(p)$ to $\mathbb{C}^{r}\otimes\ell^{2}(\Gamma\setminus\{e\})$. For $z\in\mathbb{C}\setminus\sigma(\lambda^{(e)}(p))$, we denote by
@@
the resolvent in the subtree of $\mathrm{Cay}(F_{d},(g_{i})_{i\in\left[2d\right]})$ emanating from $g_{i}$, and we let $Z^{z}_{i}=-\gamma^{z}_{i}p_{g_{i}}\in M_{r}(\mathbb{C})$. We first prove that for any $g\in\Gamma$ written in reduced form as $g=g_{i_{\ell}}\ldots g_{i_{1}}$, with $i_{k}\in\left[2d\right]$, we have
$R^{z}(g,e)=Z^{z}_{i_{\ell}}\ldots Z^{z}_{i_{1}}R^{z}(e,e).$ (6.18)
For $i\leq 2d$, we denote by $V_{i}\subset\Gamma$ the subset of group elements $h\in\Gamma$ written in reduced form $h=g_{j_{k}}\ldots g_{j_{2}}g_{j_{1}}$ with $j_{1}=i$ and $k\in\mathbb{N}$. A sequence $(j_{1},\ldots,j_{k})$ such that $g=g_{j_{k}}\ldots g_{j_{1}}$ determines a walk of length $k$ in $\Gamma$ starting from $e$ and ending at $g$. Since $\Gamma$ is the free group and the $g_{i}$’s are the free generators, we may decompose this walk at the last passage at $e$, $g_{i_{1}}$, $g_{i_{2}}g_{i_{1}}$, $\ldots$, $g_{i_{\ell}}\ldots g_{i_{1}}=g$. This induces a decomposition of the sequence $(j_{1},\ldots,j_{k})$ into $\ell+1$ sequences $s_{0},\ldots,s_{\ell}$ (possibly empty). For any $0\leq t\leq\ell$, the sequence $s_{t}$ is a walk from $g_{i_{t}}$ to $g_{i_{t}}$ in $V_{i_{t}}$ and $(j_{1},\ldots,j_{k})=(s_{0},i_{1},s_{1},\ldots,i_{\ell},s_{\ell})$. Writing $s_{t}$ as $s_{t}=(j_{t,1},\ldots,j_{t,n_{t}})$, we thus have (setting just here $p_{i}=p_{g_{i}}$)
$\sum_{k}\frac{\lambda(p)^{k}(g,e)}{z^{k+1}}=\sum_{k}\sum_{\begin{subarray}{c}(j_{1},\dots,j_{k})\\ g=g_{j_{k}}\cdots g_{j_{1}}\end{subarray}}\frac{p_{j_{k}}\cdots p_{j_{1}}}{z^{k}}=\sum_{s_{0},\dots,s_{\ell}}\frac{p_{j_{\ell,n_{\ell}}}\cdots p_{j_{\ell,1}}}{z^{n_{\ell}}}\frac{p_{i_{\ell}}}{z}\cdots\frac{p_{j_{1,n_{1}}}\cdots p_{j_{1,1}}}{z^{n_{1}}}\frac{p_{i_{1}}}{z}\frac{p_{j_{0,n_{0}}}\cdots p_{j_{0,1}}}{z^{n_{0}+1}}$
where the sum is over all sequences $(s_{0},\ldots,s_{\ell})$ of non-negative lengths $(n_{0},\ldots,n_{\ell})$ as above. Observe now that for $z$ with large enough modulus,
$R^{z}(e,e)=-\sum_{n_{0}\geq 0}\frac{\lambda(p)^{n_{0}}(e,e)}{z^{n_{0}+1}}=-\sum_{n_{0}\geq 0}\sum_{s_{0}}\frac{p_{j_{0,n_{0}}}\ldots p_{j_{0,1}}}{z^{n_{0}+1}}$ (6.19)
where the sum runs over all sequences $s_{0}=(j_{0,1},\ldots,j_{0,n_{0}})$ of length $n_{0}$ from $e$ to $e$. Similarly,
$Z_{t}^{z}=-\gamma_{t}^{z}p_{i_{t}}=\sum_{n_{t}\geq 0}\sum_{s_{t}}\frac{p_{j_{t,n_{t}}}\ldots p_{j_{t,1}}}{z^{n_{t}}}\frac{p_{i_{t}}}{z}$
where the sum is over all sequences $s_{t}=(j_{t,1},\ldots,j_{t,n_{t}})$ of length $n_{t}$ from $g_{i_{t}}$ to $g_{i_{t}}$ in $V_{i_{t}}$. This concludes the proof of (6.18) for $z$ with large enough modulus, and thus for any $z$ in the upper half-plane since both expressions in (6.18) are analytic.
@@
$\gamma_{i}^{z}=-\Bigl(z\text{Id}_{r}+\sum_{j\neq i^{*}}p_{g_{j}}^{*}\gamma_{j}^{z}p_{g_{j}}\Bigr)^{-1}.$ (6.20)
@@
Let us now specify the relations (6.18) and (6.20) to the case of $p$ of the form (6.17). In this case, $p_{g_{i}}=a_{i}E_{v_{i}u_{i}}$ for all $i\in\left[2d\right]$, where for $i\in\left[d\right]$, we have set $a_{i^{*}}=\bar{a}_{i}$ and $(u_{i^{*}},v_{i^{*}})=(v_{i},u_{i})$. In this case the relation (6.20) takes a simpler form: indeed, the matrix $p_{g_{j}}^{*}\gamma_{j}^{z}p_{g_{j}}$ is diagonal, and has only one non-vanishing coefficient, in the entry $(u_{j},u_{j})$. Let $\hat{\gamma}_{i}^{z}=(\gamma_{i}^{z})_{v_{i}v_{i}}$, then (6.20) implies the scalar equation